I'm having alot of trouble with this problem. Can you help?
Solve the equation.
square root of (2a+1)=a-1

Answer by mukhopadhyay(490)

square root of (2a+1)=a-1
=>(2a+1) = (a-1)^2 (by squaring both sides of the equation)
=>(2a+1) = a^2-2a+1
=>a^2-2a-2a+1-1 = 0
=>a^2-4a = 0
=>a(a-4) = 0
=>a = 0 or a = 4
Answer: a=0 or a=4